# Signed Networks: theory, methods, and applications

### v1

@@ 4.4.2 Spectral methods
Spectral methods for community detection infer clusters by analysing the eigenvectors of graph-based matrices, usually the Laplacian. The idea is to use the eigenvectors associated with the smallest (nontrivial) eigenvalues of the Laplacian to embed the network’s nodes in a low-dimensional space, and then identify communities as compact groups of points in that space [vonLuxburg2007a].
@@
Building on this principle, Kunegis et al. [kunegisSpectralAnalysisSigned2010, kunegisApplicationsStructuralBalance2014] proposed a natural extension of spectral methods to signed networks, arguing that it preserves the mathematical and physical interpretations that make spectral clustering effective in the unsigned case. Formally, they conceptualised communities as partitions that minimise the signed ratio cut,
@@
Building on these results, later work proposed alternative cut functions and Laplacians to address some of the limitations of Kunegis’s formulation. Chiang et al. [chiangScalableClusteringSigned2012] showed that the signed Laplacian works well for bipartitions but does not extend naturally to k-groups clustering. They introduced the Balance Ratio Cut, an objective designed to handle multiple communities more effectively, and demonstrated its advantages both analytically and numerically. More recently, Cucuringu et al. [cucuringuSPONGEGeneralizedEigenproblem2019] developed the SPONGE algorithm, which generalises the spectral approach by combining the Laplacians of the positive and negative subgraphs alongside extra tunable parameters. Their results show that SPONGE performs especially well in sparse graphs and when the number of communities is large, outperforming earlier spectral methods.
Other authors have investigated alternative matrix representations beyond the opposing Laplacian. One line of work uses the eigenvectors of the adjacency matrix [bonchiDiscoveringPolarizedCommunities2019] or of the Hashimoto (non-backtracking) matrix [zhongEfficientAlgorithmBased2020] to capture community structure directly from connectivity patterns. A related idea is to use the Gremban expansion, which embeds a signed graph into an unsigned one with twice as many nodes, so that standard spectral techniques apply without modification [fox2017numerical, foxInvestigationSpectralClustering2018].
@@ 4.4.3 Statistical inference
Most of the methods presented in previous sections had some common grounds: they relied on principles similar to structural balance and/or frustration, i.e., the aim was to find maximally balanced communities, thus with minimal frustration. A complementary approach to community detection in signed networks uses statistical inference [casellaStatisticalInference2024]. These methods assume that the network arises from a probabilistic model where edge signs depend on a latent group structure [newmanMixtureModelsExploratory2007, peixotoBayesianStochasticBlockmodeling2019], not necessarily dictated by structural balance theory. The aim is to recover the partition that best explains the observed connections, often in terms of maximum likelihood or posterior probability. Unlike quality-function-based methods, statistical models allow uncertainty estimates and can incorporate prior knowledge, especially under Bayesian frameworks [peixotoDescriptiveVsInferential2023]. They also help prevent overfitting and avoid the resolution limit. An illustration of this inferential approach to community detection is depicted in Figure 12.
The first attempt to apply statistical inference to signed networks was by Chen et al. [chenOverlappingCommunityDetection2014]. They detected communities by maximising a likelihood function under a stochastic block model, which assumed that positive links occur only within groups and negative ones only between groups. This formulation is too strict for most real networks, and as a consequence, this model is not able to detect communities in weakly balanced networks or when violations of structural balance are present. Jiang [jiangStochasticBlockModel2015] improved on Chen’s model by relaxing its assumptions. Instead of forbidding positive links between groups and negative links within groups, their model allows both types of edges to occur anywhere in the network, but with different probabilities depending on the group membership. This makes the model more flexible and capable of handling unbalanced networks.
@@
In general, the likelihood quantifies the probability of observing a given network under a particular set of model parameters. The assumption is that edges are conditionally independent given the latent variables; hence, the likelihood of the full network is written as the product of the likelihoods of individual edges. Inside the product in Eq. 8, the first term models the probability of observing a positive edge between nodes $i$ and $j$, while the second term models the probability of observing a negative edge. These probabilities depend on three components: the probability that node $i$ belongs to group $r$ as a sender ($\theta_{ri}$); the probability that node $j$ belongs to group $s$ as a receiver ($\phi_{sj}$); and the probability that nodes from groups $r$ and $s$ are connected by a positive or negative link ($\omega^{\pm}_{rs}$). Note that, unlike the hard constraint in [chenOverlappingCommunityDetection2014], here the model allows different connection probabilities for positive and negative links between any pair of groups. This gives the model greater flexibility and makes it possible to account for structural noise or weak balance; for instance, by assigning a small but nonzero probability to positive edges between antagonistic groups or to negative edges within communities. Moreover, the model generalises beyond frustration-based community detection: it can capture overlapping community structure, directed interactions (via asymmetric $\theta$ and $\phi$), and mesoscale patterns such as core–periphery or anti-balanced structures.
@@
Another approach to handling negative edges within an inferential framework was proposed by Peixoto in [peixotoNonparametricWeightedStochastic2018]. Peixoto treats edge signs as covariates (i.e., edge attributes) decoupled from the underlying network model. The network structure is modelled with a standard, degree-corrected, SBM [peixotoNonparametricBayesianInference2017, peixotoBayesianStochasticBlockmodeling2019], while the signs are accounted for independently. For real-valued signed weights, they are modeled with a Gaussian-like model; for binary signs ($\pm 1$), a Bernoulli model; and for correlations in [1,1][-1,1], the author suggests to transform the data to $(-\infty,\infty)$ using $y=2\,\mathrm{arctanh}(x)$. In all cases, this allows for avoiding overfitting and enables hierarchical inference [^22] This can be done by marginalising the weight distributions, or using instead microcanonical models that fix statistics like mean and variance. Importantly, the model does not incorporate any assumption of structural balance, which is both a strength and a limitation: on one hand, it can reveal signed group structure beyond what balance theory predicts; on the other, the inferred communities may be difficult to interpret or relate to established theoretical notions.
Other extensions of SBM-based inference include maximum-entropy approaches  [galloAssessingFrustrationRealworld2024]; models that generalise to edge weights from any exponential family [aicherLearningLatentBlock2015]; Bayesian approaches like [yangStochasticBlockmodelingVariational2017] and  [zhaoNetworkModellingVariational2018]; and EM-based variants such as [zhaoStatisticalInferenceCommunity2017, liSSBMSignedStochastic2023].
@@ 5 Dynamics
This chapter is structured as follows. Section 5.1 focuses on dynamics on networks, where the topology is fixed and only node states evolve, with emphasis on opinion dynamics models and contagion models. Section 5.2 then examines the dynamics of signed networks, in which the topology itself evolves over time and node-state dynamics are absent. We review both theoretical approaches, based on families of toy models guided by heuristic rules, and empirical studies that analyse real datasets (particularly in social contexts) to identify the mechanisms that shape network evolution. The chapter concludes with a discussion of co-evolutionary models that combine both perspectives (Section 5.3).
@@ 5.1 Dynamics on networks
In opinion dynamics, nodes are agents, and their state is their opinion on a given topic. The network edges may encode friendships, acquaintances, or connections in a digital platform. Agents repeatedly adjust their opinion based on interactions with their neighbours. Over the years, the models have incorporated various mechanisms such as cognitive biases, advertising campaigns or recommendation systems [starniniOpinionDynamicsStatistical2025]. The central question when studying these models is whether all agents eventually agree (consensus), or do they fragment into clusters with discordant opinions (polarisation). Early models of opinion dynamics were built on assimilative forces, as interactions between two agents brought their opinions closer to one another. It was not long until subsequent refinements incorporated repulsive effects [liVoterModelSigned2015], by which interactions push agents away from each other’s views (a phenomenon often referred to as the backfire effect [Nyhan2010]). Signed networks emerged as a natural framework to model this distinction: positive links promote convergence, while negative ones induce divergence or opposition. Within this framework, we distinguish two broad classes of models: diffusive dynamics and voter models.
@@ 5.1.1 Diffusive dynamics
In diffusive dynamics, opinions spread by local averaging, referring to the physical diffusion equation. These dynamics were initially introduced by French [french] and Degroot [degrootReachingConsensus1974]. Agents are endowed with opinions that they reassess at each step by taking the average of their neighbours’ opinions. Eventually consensus is reached, provided there exists a node that can reach all the others [^23] If not, a consensus is reached independently on each of the largest sub-networks to verify this property..
Altafini was the first to extend this framework to signed networks [altafiniDynamicsOpinionForming2012, altafiniConsensusProblemsNetworks2013]. In Altafini’s model, the evolution of the opinion of an agent $i$, denoted by $x_{i}$, is described by the following equation:
@@
This can be rewritten in terms of the repelling Laplacian (see Section 2.3): $\frac{dx}{dt}=-L_{r}x$. In this case, when the graph is balanced, the opinions within each faction move increasingly away from those of the opposing faction, eventually diverging. In unbalanced networks, the system may converge to consensus or diverge, depending on the network topology and the proportion of negative edges [cisneros-velardePolarizationFluctuationsSigned2021, shiDynamicsSignedNetworks2019, shiEvolutionBeliefsSigned2016].
These models have been extended in a number of ways, including time-varying topologies [proskurnikovOpinionDynamicsSocial2016], non-linearities [fontanRoleFrustrationCollective2021], and graphons [frascaOpinionDynamicsSigned2024]; they have also been applied to the study of the diversity of exposure to media outlets [quattrociocchiOpinionDynamicsInteracting2014], as well as to the problem of opinion control [zhouFriedkinJohnsenModelOpinion2024]. We refer the interested reader to [proskurnikovTutorialModelingAnalysis2018, shiDynamicsSignedNetworks2019] for more exhaustive reviews of the literature on diffusive models.
@@ 5.1.2 Voter model and random walks
The voter model is one of the most widely studied approaches to opinion dynamics. In this model, agents randomly update their state by copying that of a randomly chosen neighbour. This model’s popularity can be attributed to four key reasons. First, it offers a natural framework for herding behaviour and emergent phenomena: each agent updates its state by adopting a neighbour’s state, so large‐scale phenomena emerge from simple, local imitation. Second, the probability distribution of each agent’s opinion is a linear combination of their neighbours’ distributions, enabling analytical tractability. Third, the model has no control parameters, meaning that, apart from stochasticity, its behaviour is determined entirely by the network structure and the initial state. Finally, the model can reproduce macroscopical features of election outcomes, and retrieve individual opinions of online social media users (see [vendevilleVoterModelCan2025] and references therein).
Adaptations of the model to signed networks initially focused on the “anti‑voter” case, where all edges are negative [matloffErgodicityConditionsDissonant1977, matloffDissonantVotingModel1980, huberStationaryDistributionAntivoter2004], and then extended to systems containing both positive and negative links [saadaModeleVotantMilieu1995, gantertVoterModelAntivoter2005]. In the signed voter model, the elementary update step is modified to account for both assimilative (positive) and repulsive (negative) interactions. Assuming two possible opinions +1+1 and 1-1, the dynamics unfold in successive steps as follows. A random node $i$ is selected, and a random neighbour $j$ of $i$ is selected. Then, $i$ adopts $j$’s opinion if the $(i,j)$ edge is positive; otherwise, it adopts the opposite opinion. Let $x(t)$ be the vector whose ii‑th entry is the probability that agent $i$ is in state +1+1 at time $t$. The time evolution of $x(t)$ is governed by the equation:
@@
where $K$ is the underlying degree matrix. The first term in the equation captures the copying mechanism through positive edges, while the second term captures disagreement through negative edges. Because the dynamics are linear in $x$, Eq. 11 can be solved analytically and is formally equivalent to the evolution equation of a signed random walk [liInfluenceDiffusionDynamics2013, liVoterModelSigned2015] [^25] Recent work [babul2025strong] distinguishes strong and weak random walks on signed networks. Weak walks can handle structures with more than two antagonistic factions.. This equivalence allows the use of existing results on signed random walks [tianSpreadingStructuralBalance2024] to describe the opinion distribution. In particular, the long‑term behaviour depends exclusively on the network’s structural balance: for a balanced graph, the system converges to a static, perfectly polarized state; for an antibalanced graph [hararyStructuralDuality1957], it also achieves perfect polarization but with opinions flipping at each time step, so that the level of polarization is preserved; and for a strictly unbalanced graph, the system settles into an equilibrium in which each node holds opinion 0 or 1 with probability $1/2$, irrespective of the initial conditions. Figure 13 illustrates these different asymptotic behaviours. Building on this, subsequent work has extended the voter model to incorporate more complex interaction patterns, like directed, weighted signed graphs [liInfluenceDiffusionDynamics2013, liVoterModelSigned2015]. Other extensions include non-linear interactions [louThresholdQvoterModel2021, krawieckiVoterModelIndependence2024], mechanisms of free will [krawieckiVoterModelIndependence2024], and settings with more than two possible opinions [linOpinionPropagationSigned2019].
Figure 13: Dynamics and asymptotic behaviour of the Voter Model with synchronous updates on signed networks [liInfluenceDiffusionDynamics2013, liVoterModelSigned2015]. Node colours designate opinions in $\{-1,+1\}$. There are two factions, distinguished by full and dashed lines. We show the networks with equilibrium opinions as well as dynamics over time.
@@ 5.1.3 Compartmental models
Figure 14: Evolution of the SIS (Susceptible-Infected-Susceptible), IC (Independent Cascade), and LT (Linear Threshold) models. Contagion-like models usually include two opposing infected states (represented here in green and pink) and one susceptible state (in grey). Transmission along a positive edge preserves the infected state of the source node, while transmission along a negative edge flips it to the opposite infected state. Black arrows indicate the time direction, with multiple time-steps between the left and right panels for all models. LT panel adapted from [srivastavaSocialInfluenceComputation2015]
@@
In epidemic spreading models, populations are divided into health state compartments, typically Susceptible (not yet infected), Infected (currently spreading the disease), and Recovered (immune after infection). Compartmental models describe how agents move between discrete states, such as susceptible, infected, or recovered, through probabilistic transition rates that capture continuous-time spreading and recovery. One of the first applications of this framework to signed networks was proposed by [fanAnalysisOpinionSpreading2012], who extended the Susceptible-Infected-Recovered model (SIR) model to homogeneous signed networks with a varying fraction of negative edges. Susceptible ($S$) individuals hold the neutral state and can be infected by either type 1 ($I_{1}$) or type 2 ($I_{2}$) infectious individuals. When the link SIS-I is positive, the susceptible node will adopt the same state as the infected one with probability $\beta_{1}$; when it is negative, the node adopts the opposite state with probability $\beta_{2}$. Both types of infected state may transition to the recovered state at a rate $\gamma$. The authors find a threshold in the number of negative links beyond which contagion slows down and the system polarises, while networks dominated by positive links converge rapidly to a consensus.
@@
Another approach sees negative edges not as spreading an opposite state but as changing how easily a process, like adopting a new technique, spreads through the network. For example, in [liDynamicsEpidemicSpreading2021], each edge in the signed network carries its own infection rate: a positive edge transmits with a rate proportional to the fraction of balanced triads in which it participates, while a negative edge transmits with a rate proportional to the fraction of unbalanced triads it belongs to. Thus, contagion depends not only on the sign of a link but also on its local structural balance.
@@
Signed compartmental models have also been extended to multiplex and higher-order settings. In multiplex networks, studies such as [fengHeterogeneousPropagationProcesses2024, SheOnANetworkedSIS2022] couple an unsigned physical-contact layer, where SIS contagion occurs, with a signed awareness layer, where information about the disease circulates. When all links in the awareness layer are positive, individuals reinforce each other’s caution, which raises the epidemic threshold and reduces prevalence. When the awareness layer includes negative links, however, some individuals oppose their neighbours’ warnings, which weakens collective responses and traps the system in polarised states. These divisions reduce perceived risk and allow the epidemic to persist, even when awareness is widespread. In higher-order contagion models, contagion acts through both single links and group interactions. When all links are positive, infections within a group strengthen one another and favour collective activation. The presence of negative links changes this logic: negative links create feedback that can oppose, or under certain conditions, reinforce the spread. [kemmeterComplexContagionSocial2024] finds that all-negative triads slow contagion at low infection rates but accelerate it once the process becomes strong enough. Fully positive triads make bistability possible and keep the infection active, while configurations with two negative links block propagation. [maSocialContagionEmotional2024] reaches a similar conclusion: as antagonism becomes more common, collective adoption weakens, the epidemic threshold increases, and the transition to the disease-free state shifts from discontinuous to continuous.

### v2

@@ 4.4.2 Spectral methods
Spectral methods for community detection infer clusters by analysing the eigenvectors of graph-based matrices, usually the Laplacian. The idea is to use the eigenvectors associated with the smallest (nontrivial) eigenvalues of the Laplacian to embed the network’s nodes in a low-dimensional space, and then identify communities as compact groups of points in that space [115].
@@
Building on this principle, Kunegis et al. [224, 225] proposed a natural extension of spectral methods to signed networks, arguing that it preserves the mathematical and physical interpretations that make spectral clustering effective in the unsigned case. Formally, they conceptualised communities as partitions that minimise the signed ratio cut,
@@
Building on these results, later work proposed alternative cut functions and Laplacians to address some of the limitations of Kunegis’s formulation. Chiang et al. [69] showed that the signed Laplacian works well for bipartitions but does not extend naturally to k-groups clustering. They introduced the Balance Ratio Cut, an objective designed to handle multiple communities more effectively, and demonstrated its advantages both analytically and numerically. More recently, Cucuringu et al. [77] developed the SPONGE algorithm, which generalises the spectral approach by combining the Laplacians of the positive and negative subgraphs alongside extra tunable parameters. Their results show that SPONGE performs especially well in sparse graphs and when the number of communities is large, outperforming earlier spectral methods.
Other authors have investigated alternative matrix representations beyond the opposing Laplacian. One line of work uses the eigenvectors of the adjacency matrix [46] or of the Hashimoto (non-backtracking) matrix [419] to capture community structure directly from connectivity patterns. A related idea is to use the Gremban expansion, which embeds a signed graph into an unsigned one with twice as many nodes, so that standard spectral techniques apply without modification [127, 128].
@@ 4.4.3 Statistical inference
Most of the methods presented in previous sections had some common grounds: they relied on principles similar to structural balance and/or frustration, i.e., the aim was to find maximally balanced communities, thus with minimal frustration. A complementary approach to community detection in signed networks uses statistical inference [57]. These methods assume that the network arises from a probabilistic model where edge signs depend on a latent group structure [294, 309], not necessarily dictated by structural balance theory. The aim is to recover the partition that best explains the observed connections, often in terms of maximum likelihood or posterior probability. Unlike quality-function-based methods, statistical models allow uncertainty estimates and can incorporate prior knowledge, especially under Bayesian frameworks [310]. They also help prevent overfitting and avoid the resolution limit. An illustration of this inferential approach to community detection is depicted in Figure 12.
The first attempt to apply statistical inference to signed networks was by Chen et al. [66]. They detected communities by maximising a likelihood function under a stochastic block model, which assumed that positive links occur only within groups and negative ones only between groups. This formulation is too strict for most real networks, and as a consequence, this model is not able to detect communities in weakly balanced networks or when violations of structural balance are present. Jiang [202] improved on Chen’s model by relaxing its assumptions. Instead of forbidding positive links between groups and negative links within groups, their model allows both types of edges to occur anywhere in the network, but with different probabilities depending on the group membership. This makes the model more flexible and capable of handling unbalanced networks.
@@
In general, the likelihood quantifies the probability of observing a given network under a particular set of model parameters. The assumption is that edges are conditionally independent given the latent variables; hence, the likelihood of the full network is written as the product of the likelihoods of individual edges. Inside the product in Eq. 8, the first term models the probability of observing a positive edge between nodes $i$ and $j$, while the second term models the probability of observing a negative edge. These probabilities depend on three components: the probability that node $i$ belongs to group $r$ as a sender ($\theta_{ri}$); the probability that node $j$ belongs to group $s$ as a receiver ($\phi_{sj}$); and the probability that nodes from groups $r$ and $s$ are connected by a positive or negative link ($\omega^{\pm}_{rs}$). Note that, unlike the hard constraint in [66], here the model allows different connection probabilities for positive and negative links between any pair of groups. This gives the model greater flexibility and makes it possible to account for structural noise or weak balance; for instance, by assigning a small but nonzero probability to positive edges between antagonistic groups or to negative edges within communities. Moreover, the model generalises beyond frustration-based community detection: it can capture overlapping community structure, directed interactions (via asymmetric $\theta$ and $\phi$), and mesoscale patterns such as core–periphery or anti-balanced structures.
@@
Another approach to handling negative edges within an inferential framework was proposed by Peixoto in [308]. Peixoto treats edge signs as covariates (i.e., edge attributes) decoupled from the underlying network model. The network structure is modelled with a standard, degree-corrected, SBM [307, 309], while the signs are accounted for independently. For real-valued signed weights, they are modeled with a Gaussian-like model; for binary signs ($\pm 1$), a Bernoulli model; and for correlations in [1,1][-1,1], the author suggests to transform the data to $(-\infty,\infty)$ using $y=2\,\mathrm{arctanh}(x)$. In all cases, this allows for avoiding overfitting and enables hierarchical inference [^22] This can be done by marginalising the weight distributions, or using instead microcanonical models that fix statistics like mean and variance. Importantly, the model does not incorporate any assumption of structural balance. This is both a strength and a limitation. On one hand, it can reveal signed mesoscale structures beyond what balance theory predicts. Examples include core-periphery configurations and assortative structures where both positive and negative links are concentrated within the same groups. According to [415], these unbalanced patterns are prevalent in real-world networks, even those that appear balanced at the micro-scale level of triangles. On the other hand, the inferred communities may be difficult to interpret or relate to established theoretical notions, as explored in [150].
Other extensions of SBM-based inference include maximum-entropy approaches  [139, 89]; models that generalise to edge weights from any exponential family [3]; Bayesian approaches like [403] and  [417]; and EM-based variants such as [418, 238].
@@ 5 Dynamics
This chapter is structured as follows. Section 5.1 focuses on dynamics on networks, where the topology is fixed, and only node states evolve, with emphasis on opinion dynamics models and contagion models. Section 5.2 then examines the dynamics of signed networks, in which the topology itself evolves over time and node-state dynamics are absent. We review both theoretical approaches, based on families of toy models guided by heuristic rules, and empirical studies that analyse real datasets (particularly in social contexts) to identify the mechanisms that shape network evolution. The chapter concludes with a discussion of co-evolutionary models that combine both perspectives (Section 5.3).
@@ 5.1 Dynamics on networks
In opinion dynamics, nodes are agents, and their state is their opinion on a given topic. The network edges may encode friendships, acquaintances, or connections in a digital platform. Agents repeatedly adjust their opinion based on interactions with their neighbours. Over the years, the models have incorporated various mechanisms such as cognitive biases, advertising campaigns, or recommendation systems [367]. The central question when studying these models is whether all agents eventually agree (consensus), or whether they fragment into clusters with discordant opinions (polarisation). Early models of opinion dynamics were built on assimilative forces, as interactions between two agents brought their opinions closer to one another. It was not long until subsequent refinements incorporated repulsive effects [240], by which interactions push agents away from each other’s views (a phenomenon often referred to as the backfire effect [299]). Signed networks emerged as a natural framework to model this distinction: positive links promote convergence, while negative ones induce divergence or opposition. Within this framework, we distinguish two broad classes of models: diffusive dynamics and voter models.
@@ 5.1.1 Diffusive dynamics
In diffusive dynamics, opinions spread by local averaging, referring to the physical diffusion equation. These dynamics were initially introduced by French [134] and Degroot [81]. Agents are endowed with opinions that they reassess at each step by taking the average of their neighbours’ opinions. Eventually, a consensus is reached, provided there exists a node that can reach all the others [^23] If not, a consensus is reached independently on each of the largest sub-networks to verify this property..
Altafini was the first to extend this framework to signed networks [10, 9]. In Altafini’s model, the evolution of the opinion of an agent $i$, denoted by $x_{i}$, is described by the following equation:
@@
This can be rewritten in terms of the repelling Laplacian (see Section 2.3): $\frac{dx}{dt}=-L_{r}x$. In this case, when the graph is balanced, the opinions within each faction move increasingly away from those of the opposing faction, eventually diverging. In unbalanced networks, the system may converge to consensus or diverge, depending on the network topology and the proportion of negative edges [74, 353, 354].
These models have been extended in a number of ways, including time-varying topologies [319], non-linearities [121], and graphons [318]; they have also been applied to the study of the diversity of exposure to media outlets [321], as well as to the problem of opinion control [420]. We refer the interested reader to [320, 353] for more exhaustive reviews of the literature on diffusive models.
@@ 5.1.2 Voter model and random walks
The voter model is one of the most widely studied approaches to opinion dynamics. In this model, agents randomly update their state by copying that of a randomly chosen neighbour. This model’s popularity can be attributed to four key reasons. First, it offers a natural framework for herding behaviour and emergent phenomena: each agent updates its state by adopting a neighbour’s state, so large‐scale phenomena emerge from simple, local imitation. Second, the probability distribution of each agent’s opinion is a linear combination of their neighbours’ distributions, enabling analytical tractability. Third, the model has no control parameters, meaning that, apart from stochasticity, its behaviour is determined entirely by the network structure and the initial state. Finally, the model can reproduce macroscopical features of election outcomes, and retrieve individual opinions of online social media users (see [392] and references therein).
Adaptations of the model to signed networks initially focused on the “anti‑voter” case, where all edges are negative [265, 266, 193], and then extended to systems containing both positive and negative links [333, 141]. In the signed voter model, the elementary update step is modified to account for both assimilative (positive) and repulsive (negative) interactions. Assuming two possible opinions +1+1 and 1-1, the dynamics unfold in successive steps as follows. A random node $i$ is selected, and a random neighbour $j$ of $i$ is selected. Then, $i$ adopts $j$’s opinion if the $(i,j)$ edge is positive; otherwise, it adopts the opposite opinion. Let $x(t)$ be the vector whose ii‑th entry is the probability that agent $i$ is in state +1+1 at time $t$. The time evolution of $x(t)$ is governed by the equation:
@@
where $K$ is the underlying degree matrix. The first term in the equation captures the copying mechanism through positive edges, while the second term captures disagreement through negative edges. Because the dynamics are linear in $x$, Eq. 11 can be solved analytically and is formally equivalent to the evolution equation of a signed random walk [239, 240] [^25] Recent work [29] distinguishes strong and weak random walks on signed networks. Weak walks can handle structures with more than two antagonistic factions.. This equivalence allows the use of existing results on signed random walks [383] to describe the opinion distribution. In particular, the long‑term behaviour depends exclusively on the network’s structural balance: for a balanced graph, the system converges to a static, perfectly polarized state; for an antibalanced graph [170], it also achieves perfect polarization but with opinions flipping at each time step, so that the level of polarization is preserved; and for a strictly unbalanced graph, the system settles into an equilibrium in which each node holds opinion 0 or 1 with probability $1/2$, irrespective of the initial conditions. Figure 13 illustrates these different asymptotic behaviours. Building on this, subsequent work has extended the voter model to incorporate more complex interaction patterns, like directed, weighted signed graphs [239, 240]. Other extensions include non-linear interactions [249, 217], mechanisms of free will [217], and settings with more than two possible opinions [241]. The similar but deterministic majority-rule model has also received some attention [72, 88].
Figure 13: Dynamics and asymptotic behaviour of the Voter Model with synchronous updates on signed networks [239, 240]. Node colours designate opinions in $\{-1,+1\}$. There are two factions, distinguished by full and dashed lines. We show the networks with equilibrium opinions as well as dynamics over time.
@@ 5.1.3 Compartmental models
Figure 14: Evolution of the SIS (Susceptible-Infected-Susceptible), IC (Independent Cascade), and LT (Linear Threshold) models. Contagion-like models usually include two opposing infected states (represented here in green and pink) and one susceptible state (in grey). Transmission along a positive edge preserves the infected state of the source node, while transmission along a negative edge flips it to the opposite infected state. Black arrows indicate the time direction, with multiple time-steps between the left and right panels for all models. LT panel adapted from [365]
@@
In epidemic spreading models, populations are divided into health state compartments, typically Susceptible (not yet infected), Infected (currently spreading the disease), and Recovered (immune after infection). Compartmental models describe how agents move between discrete states, such as susceptible, infected, or recovered, through probabilistic transition rates that capture continuous-time spreading and recovery. One of the first applications of this framework to signed networks was proposed by [116], who extended the Susceptible-Infected-Recovered model (SIR) model to homogeneous signed networks with a varying fraction of negative edges. Susceptible ($S$) individuals hold the neutral state and can be infected by either type 1 ($I_{1}$) or type 2 ($I_{2}$) infectious individuals. When the link SIS-I is positive, the susceptible node will adopt the same state as the infected one with probability $\beta_{1}$; when it is negative, the node adopts the opposite state with probability $\beta_{2}$. Both types of infected state may transition to the recovered state at a rate $\gamma$. The authors find a threshold in the number of negative links beyond which contagion slows down, and the system polarises, while networks dominated by positive links converge rapidly to a consensus.
@@
Another approach sees negative edges not as spreading an opposite state but as changing how easily a process, like adopting a new technique, spreads through the network. For example, in [235], each edge in the signed network carries its own infection rate: a positive edge transmits with a rate proportional to the fraction of balanced triads in which it participates, while a negative edge transmits with a rate proportional to the fraction of unbalanced triads it belongs to. Thus, contagion depends not only on the sign of a link but also on its local structural balance.
@@
Signed compartmental models have also been extended to multiplex and higher-order settings. In multiplex networks, studies such as [119, 351] couple an unsigned physical-contact layer, where SIS contagion occurs, with a signed awareness layer, where information about the disease circulates. When all links in the awareness layer are positive, individuals reinforce each other’s caution, which raises the epidemic threshold and reduces prevalence. When the awareness layer includes negative links, however, some individuals oppose their neighbours’ warnings, which weakens collective responses and traps the system in polarised states. These divisions reduce perceived risk and allow the epidemic to persist, even when awareness is widespread. In higher-order contagion models, contagion acts through both single links and group interactions. When all links are positive, infections within a group strengthen one another and favour collective activation. The presence of negative links changes this logic: negative links create feedback that can oppose, or under certain conditions, reinforce the spread. [207] finds that all-negative triads slow contagion at low infection rates but accelerate it once the process becomes strong enough. Fully positive triads make bistability possible and keep the infection active, while configurations with two negative links block propagation. [253] reaches a similar conclusion: as antagonism becomes more common, collective adoption weakens, the epidemic threshold increases, and the transition to the disease-free state shifts from discontinuous to continuous.